OpenStudy (anonymous):

-6=y+25 solve and check.

OpenStudy (anonymous):

Subtract 25 from both sides to get y = -31 -6 = (-31) + 25

OpenStudy (anonymous):

sorry im not clear with dis eqn can u xplain beta the checking aspect show full working thanks

OpenStudy (anonymous):

Plugging -31 into your equation gives you -6 = -31 + 25 -31 + 25 = -6 So -6 = -6, which validates that y = -31

OpenStudy (anonymous):

k thanks alot

OpenStudy (anonymous):

Complete nonsense, the only thing u need to to check is that u can take 25 away properly, duh!